OpenStudy (ilovepurple42043):

a cylinder has a radius of 2x + 3 and a height of 6x + 1. Which polynomial in standard form best describes the total volume of the cylinder? use the formula v=pir^2h for the volume of a cylinder. (1 point)

OpenStudy (anonymous):

just plug r=2x+3 h=6x+1 and simplify

OpenStudy (anonymous):

\[\left( a+b \right)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

cubic polynomial.